# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
We also recall that by construction $B$ with the induced metric is geodesic in $(P_{\pm}\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/K_{\pm}$ with respect to the the two $\mathsf{RCD}(0,N^{F}_{\pm}+1+N^{P}_{\pm})$-structures $(d_{\pm},\mathfrak{m}_{\pm})$. Moreover, over the open collars $(P_{\pm}\times((1/A_{\pm},2/A_{\pm}]\times F_{\pm}))/G_{\pm}$ the metric structures $(d_{\pm},\mathfrak{m}_{\pm})$ are the product structures $(d_{\mathbb{R}}\oplus d^{P}_{\pm}\oplus d^{F}_{\pm},f^{N^{F}_{\pm}}_{\pm}d\mathcal{L}\otimes d\mathfrak{m}^{P}_{\pm}\otimes d\mathfrak{m}^{F}_{\pm})$. From this we observe that induced metric structures on $B$ are of the form $((d^{P}_{\pm}\otimes d^{F}_{\pm})^{\ast},(\mathfrak{m}^{P}\pm\otimes\mathfrak{m}^{F}_{\pm})_{\ast})$ and are given by a Riemannian metric $g_{\pm}$ on $B$: since $P_{\pm}\times F_{\pm}$ with $g_{P_{\pm}}\oplus g_{F_{\pm}}$ is a Riemannian manifold inducing $(d^{P}_{\pm}\oplus d^{F}_{\pm},\mathfrak{m}^{P}_{\pm}\oplus\mathfrak{m}^{F}_{\pm})$, and the action of $G_{\pm}$ over $P_{\pm}\times F_{\pm}$ is free, we obtain quotient Riemannian metrics $g_{\pm}$ on $B$ inducing $((d^{P}_{\pm}\oplus d^{F}_{\pm})^{\ast},(\mathfrak{m}^{P}_{\pm}\otimes\mathfrak{m}^{F}_{\pm})_{\ast})$.
We consider $[-1,1]\times B$, and consider $g(t)$ a path in the set of Riemannian metrics on $B$ with $g(t)=g_{-}$ over $[-1,-1+\varepsilon]$ and $g(t)=g_{+}$ over $[1-\varepsilon,1]$. We set $C_{0}:=[-1,1]\times B$ and equip it with the metric $d_{0}$ and measure $\mathfrak{m}_{0}$ induced by the Riemannian metric $h=dt^{2}\oplus g(t)$. We observe that since $([-1,1]\times B,h)$ is a compact Riemannian manifold, then there exists $K_{0}\in\mathbb{R}$, $|K_{0}|\leq\infty$ such that $\operatorname{Ric}(h)\geq K_{0}$. This implies that $(C_{0},d_{0},\mathfrak{m}_{0})$ is an $\mathsf{RCD}(K_{0},\dim(B)+1)$-space.
We identify $(P_{-}\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ with $C_{-}:=(P_{-}\times([-1-2/A_{-},-1]\times F/(-1-2/A_{-},p)\sim(-1-2/A_{-},q))/G_{-}$ and push the $\mathsf{RCD}(0,N^{F}_{-}+1+N^{P}_{-})$-structure $(d_{-},\mathfrak{m}_{-})$ of $(P_{-}\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ to get an $\mathsf{RCD}(0,N^{F}_{-}+1+N^{P}_{-})$-structure on $C_{-}$, which we also denote by $(d_{-},\mathfrak{m}_{-})$ abusing notation. We also identify $(P_{+}\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ with $C_{+}:=(P_{+}\times([1,1+2/A_{+}]\times F/(1+2/A_{+},p)\sim(1+2/A_{+},q))/G_{+}$ and push the $\mathsf{RCD}(0,N^{F}_{+}+1+N^{P}_{+})$-structure $(d_{+},\mathfrak{m}_{+})$ of $(P_{+}\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ to get an $\mathsf{RCD}(0,n_{+}1+D)$-structure on $C_{+}$, which we also denote by $(d_{+},\mathfrak{m}_{+})$.
@@
$d_{1}(x,y)=\begin{cases}d_{-}(x,y),&x,y\in C_{-}\setminus\partial C_{-}=C_{-}\setminus B,\\ d_{0}(x,y),&x,y\in C_{0}\setminus(\{-1\}\times B),\\ \min\{d_{-}(x,a)+d_{0}((-1,a),y)\mid a\in B\},&x\in C_{-}\mbox{ and }y\in C_{0}.\\ \end{cases}$
We also set the measure $\mathfrak{n}_{1}:=(i_{-})_{\ast}(\mathfrak{m}_{-})+(i_{0})_{\ast}(\mathfrak{m}_{0})$, where $i_{-}\colon C_{-}\to X$ and $i_{0}\colon C_{0}\to X$ are the respective inclusions. Also observe that for $K_{1}=\min\{K_{0},0\}$ and $N_{1}=\max\{N^{F}_{-}+1+N^{P}_{\pm},\dim(B)+1\}$, we have that $C_{0}$, $C_{-}$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces. We consider $\delta_{1}<\min\{\varepsilon,1/A_{-}\}$ as well as the open subsets $U_{-}=C_{-}\cup([-1,-1+\delta)\times B\subset X_{1}$, and $U_{0}=(-1-\delta,1]\times B$, and observe that they are an open cover of $X_{1}$ due to the fact that the metrics $d_{-}$ and $d_{0}$ are the product metrics over the preimage of $[-1-(2/A_{-})-\delta,-1+\delta]\subset X_{1}^{\ast}$ under the quotient map $\pi\colon X_{1}\to X_{1}^{\ast}$. Moreover, since the metric $d_{1}$ and measure $\mathfrak{m}_{1}$ are the product over $[-1-\delta,-1+\delta]$, we have that the spaces $(\bar{U}_{-},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{-}})$, $(\bar{U}_{0},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{0}})$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces by construction. Thus by Theorem 2.11, we conclude that $X_{1}$ is an $\mathsf{RCD}(K_{1},N_{1})$-space.
In analogous fashion we can prove that the spaces
@@
We now prove Theorem B.
Proof of Theorem B.
We set $F_{\pm}$ to be the homogeneous Riemannian manifolds $G_{\pm}/H$ with the $G$-invariant Riemannian metrics of positive Ricci curvature. Then they are $\mathsf{RCD}((n_{\pm}-1)K_{\pm},n_{\pm})$-spaces, for $n_{\pm}=\dim(F_{\pm})$ and some $K_{\pm}>0$ fixed. Observe that in the case when either $n_{+}=1$ or $n_{-}=1$, then by taking $\lambda_{\pm}\geq\operatorname{diam}(F_{\pm})/(\pi/\sqrt{K_{\pm}})$, we have that by rescaling $F_{\pm}$ by $\lambda_{\pm}$, the rescaled space $(F_{\pm},\lambda_{\pm}d_{F_{\pm}},(\lambda_{\pm})^{n_{\pm}}\mathfrak{m}_{F_{\pm}})$ is an $\mathsf{RCD}((n_{\pm}-1)K_{\pm}/\lambda_{\pm},n_{\pm})$-space which satisfies $\operatorname{diam}(F_{\pm},\lambda_{\pm}d_{F_{\pm}})\leq\pi\lambda_{\pm}^{2}/\sqrt{K_{\pm}}$. Thus without loss of generality, we may assume that if either $n_{+}=1$ or $n_{-}=1$, we have $\operatorname{diam}(F_{\pm})\leq\pi/\sqrt{K_{\pm}}$.
@@
$\Big{(}(G\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/G_{\pm},(d_{G}\otimes d_{A_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes d\mathfrak{m}_{A_{\pm}})_{\ast}\Big{)},$
[…]
We showed in Proposition 5.11 that over $(1/A_{\pm},2/A_{\pm}]\times F_{\pm}\subset[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}$ we have $d_{A_{\pm}}=d_{\mathbb{R}}\oplus d_{F_{\pm}}$. Since the action of $G_{\pm}$ on the $[0,2/A\pm]$ factor is free, we see that the image of $G\times[1/A_{\pm},2/A_{\pm}]\times F_{\pm}$ under the quotient map $G\times([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})\to(G\times([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}))/\Delta$ is isometric to
@@
We also recall that by construction the space $B$ with the induced metric is geodesic in $(G\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/G_{\pm}$ with respect to the the two $\mathsf{RCD}(0,n_{\pm}+1+D)$-structures $(d_{\pm},\mathfrak{m}_{\pm})$. Moreover, over the open collars $(G((1/A_{\pm},2/A_{\pm}]\times F_{\pm}))/G_{\pm}$ the metric structures $(d_{\pm},\mathfrak{m}_{\pm})$ are the product structures $(d_{\mathbb{R}}\oplus d_{G}\oplus d_{F_{\pm}},f^{n_{\pm}}_{\pm}d\mathcal{L}\otimes d\mathfrak{m}_{G}\otimes d\mathfrak{m}_{F_{\pm}})$. From this we observe that induced metric structures on $B$ are of the form $((d_{G}\otimes d_{F_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes\mathfrak{m}_{F_{\pm}})_{\ast})$ and are given by a Riemannian metric $g_{\pm}$ on $B$: since $G\times F_{\pm}$ with $g_{G}\oplus g_{F_{\pm}}$ is a Riemannian manifold inducing $(d_{G}\oplus d_{F_{\pm}},\mathfrak{m}_{G}\oplus\mathfrak{m}_{F_{\pm}})$, and the action of $G_{\pm}$ over $G\times F_{\pm}$ is free, we obtain quotient Riemannian metrics $g_{\pm}$ on $B$ inducing $((d_{G}\oplus d_{F_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes\mathfrak{m}_{F})_{\ast})$.
We consider $[-1,1]\times B$, and consider $g(t)$ a path in the set of Riemannian metrics on $B$ with $g(t)=g_{-}$ over $[-1,-1+\varepsilon]$ and $g(t)=g_{+}$ over $[1-\varepsilon,1]$. We set $C_{0}:=[-1,1]\times B$ and equip it with the metric $d_{0}$ and measure $\mathfrak{m}_{0}$ induced by the Riemannian metric $h=dt^{2}\oplus g(t)$. We observe that since $([-1,1]\times B,h)$ is a compact Riemannian manifold, then there exists $K_{0}\in\mathbb{R}$, $|K_{0}|\leq\infty$ such that $\operatorname{Ric}(h)\geq K_{0}$. This implies that $(C_{0},d_{0},\mathfrak{m}_{0})$ is an $\mathsf{RCD}(K_{0},\dim(B)+1)$-space.
We identify $[0,2/A_{-}]\times_{f_{-}}F_{-}$ with $C_{-}:=(G\times([-1-2/A_{-},-1]\times F/(-1-2/A_{-},p)\sim(-1-2/A_{-},q)))/G_{-}$ and push the $\mathsf{RCD}(0,n_{-}+1+D)$-structure $(d_{-},\mathfrak{m}_{-})$ of $(G\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ to get an $\mathsf{RCD}(0,n_{-}+1+D)$-structure on $C_{-}$, which we also denote by $(d_{-},\mathfrak{m}_{-})$ abusing notation. We also identify $G\times([0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ with $C_{+}:=(G\times([1,1+2/A_{+}]\times F/(1+2/A_{+},p)\sim(1+2/A_{+},q)))/G_{+}$ and push the $\mathsf{RCD}(0,n_{+}1+D)$-structure $(d_{+},\mathfrak{m}_{+})$ of $(G\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ to get an $\mathsf{RCD}(0,n_{+}1+D)$-structure on $C_{+}$, which we also denote by $(d_{+},\mathfrak{m}_{+})$.
@@
$d_{1}(x,y)=\begin{cases}d_{-}(x,y),&x,y\in C_{-}\setminus\partial C_{-}=C_{-}\setminus B,\\ d_{0}(x,y),&x,y\in C_{0}\setminus(\{-1\}\times B),\\ \min\{d_{-}(x,a)+d_{0}((-1,a),y)\mid a\in B\},&x\in C_{-}\mbox{ and }y\in C_{0}.\\ \end{cases}$
We also set the measure $\mathfrak{n}_{1}:=(i_{-})_{\ast}(\mathfrak{m}_{-})+(i_{0})_{\ast}(\mathfrak{m}_{0})$, where $i_{-}\colon C_{-}\to X$ and $i_{0}\colon C_{0}\to X$ are the respective inclusions. Also observe that for $K_{1}=\min\{K_{0},0\}$ and $N_{1}=\max\{n_{-}+1+D,\dim(B)+1\}$, we have that $C_{0}$, $C_{-}$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces. We consider $\delta_{1}<\min\{\varepsilon,1/A_{-}\}$ as well as the open subsets $U_{-}=C_{-}\cup([-1,-1+\delta)\times B\subset X_{1}$, and $U_{0}=(-1-\delta,1]\times B$, and observe that they are an open cover of $X_{1}$ due to the fact that the metrics $d_{-}$ and $d_{0}$ are the product metrics over the preimage of $[-1-(2/A_{-})-\delta,-1+\delta]\subset X_{1}^{\ast}$ under the quotient map $\pi\colon X_{1}\to X_{1}^{\ast}$. Moreover, since the metric $d_{1}$ and measure $\mathfrak{m}_{1}$ are the product over $[-1-\delta,-1+\delta]$, we have that the spaces $(\bar{U}_{-},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{-}})$, $(\bar{U}_{0},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{0}})$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces by construction. Thus by Theorem 2.11, we conclude that $X_{1}$ is an $\mathsf{RCD}(K_{1},N_{1})$-space. Observe that since the action of $G$ is transitive on $B$, we have that
(5) $\min\{d_{-}(x,a)$ $+d_{0}((-1,a),y)\mid a\in B\}$
(6) $=\min\{d_{-}(x,g^{-1}a)+d_{0}((-1,g^{-1}a),y)\mid g^{-1}a\in B\}$
(7) $=\min\{d_{-}(x,\tilde{a}])+d_{0}((-1,\tilde{a}),y)\mid\tilde{w}\in B\}.$
@@
Let $x_{0}\in X$ be contained in a principal orbit, and take $k=\mathrm{dim}(G(x_{0}))$. Recall that the orbit $G(x_{0})$ is homeomorphic to the homogeneous space $G/G_{x_{0}}$. By [14, Chapter III, Section IV 54], there exists a $G$-invariant Riemannian metric $g_{0}$ on $G/G_{x_{0}}$ such that $G\subset\mathrm{Isom}(G/G_{x_{0}},g_{0})$. Hence, due to [63] and [55, Chapter II, Section 3, Theorem 3.1], we have the following dimension estimate
@@
We begin by pointing out that, if $X^{*}=\mathbf{S}^{1}$, then $X$ is homeomorphic to a $3$-manifold by item (c) of Theorem A, and therefore the statement holds true. Thus, in the following we assume that $X^{*}=[-1,1]$.
@@
Assume first that $G=T^{3}$, and observe that all principal isotropy groups are the same for this action, since they are conjugated to each other (see [27, Theorem 4.7 and proof of Corollary 4.9]) and $G$ is abelian. Moreover, as $G$ acts by cohomogeneity one on $X$ we have a group diagram of the form:
@@
We recall from [27, Corollary 10] that $\pi^{-1}(-1,1)$ is homeomorphic to $(-1,1)\times G_{x}$, where $x\in\pi^{-1}(-1,1)$.
@@
Now from Theorem 2.22 we have that $X$ is non-collapsed, and we consider $x_{\pm}\in X$ such that $x_{\pm}^{\ast}=\pm 1$. We study the tubular neighborhoods of the orbit $G(x_{\pm})$. We denote by $m_{\pm}=\mathrm{dim}(G(x_{\pm}))$, and observe that by our hypothesis $N=3$. Due to Theorem 2.29, we have that a tubular neighborhood of the orbit $G(x_{\pm})$ is homeomorphic to $G\times_{G_{\pm}}S_{\pm}$, where $S_{\pm}$ is the slice through $x_{\pm}$.

### v3

@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
We also recall that by construction $B$ with the induced metric is geodesic in $(P_{\pm}\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/G_{\pm}$ with respect to the the two $\mathsf{RCD}(\bar{K}_{\pm},N^{F}_{\pm}+1+N^{P}_{\pm})$-structures $(d_{\pm},\mathfrak{m}_{\pm})$. Moreover, over the open collars $(P_{\pm}\times((1/A_{\pm},2/A_{\pm}]\times F_{\pm}))/G_{\pm}$ the metric structures $(d_{\pm},\mathfrak{m}_{\pm})$ are the product structures $(d_{\mathbb{R}}\oplus d^{P}_{\pm}\oplus d^{F}_{\pm},f^{N^{F}_{\pm}}_{\pm}d\mathcal{L}\otimes d\mathfrak{m}^{P}_{\pm}\otimes d\mathfrak{m}^{F}_{\pm})$. From this we observe that induced metric structures on $B$ are of the form $((d^{P}_{\pm}\otimes d^{F}_{\pm})^{\ast},(\mathfrak{m}^{P}\pm\otimes\mathfrak{m}^{F}_{\pm})_{\ast})$ and are given by a Riemannian metric $g_{\pm}$ on $B$: since $P_{\pm}\times F_{\pm}$ with $g_{P_{\pm}}\oplus g_{F_{\pm}}$ is a Riemannian manifold inducing $(d^{P}_{\pm}\oplus d^{F}_{\pm},\mathfrak{m}^{P}_{\pm}\oplus\mathfrak{m}^{F}_{\pm})$, and the action of $G_{\pm}$ over $P_{\pm}\times F_{\pm}$ is free, we obtain quotient Riemannian metrics $g_{\pm}$ on $B$ inducing $((d^{P}_{\pm}\oplus d^{F}_{\pm})^{\ast},(\mathfrak{m}^{P}_{\pm}\otimes\mathfrak{m}^{F}_{\pm})_{\ast})$.
We consider $1,1]\times B$, and consider $g(t)$ a path in the set of Riemannian metrics on $B$ with $g(t)=g_{-}$ over $1,-1+\varepsilon]$ and $g(t)=g_{+}$ over $1-\varepsilon,1]$. We set $C_{0}:=[-1,1]\times B$ and equip it with the metric $d_{0}$ and measure $\mathfrak{m}_{0}$ induced by the Riemannian metric $h=dt^{2}\oplus g(t)$. We observe that since $([-1,1]\times B,h)$ is a compact Riemannian manifold, then there exists $K_{0}\in\mathbb{R}$, $|K_{0}|\leq\infty$ such that $\operatorname{Ric}(h)\geq K_{0}$. This implies that $(C_{0},d_{0},\mathfrak{m}_{0})$ is an $\mathsf{RCD}(K_{0},\dim(B)+1)$-space.
We identify $(P_{-}\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ with $C_{-}:=(P_{-}\times([-1-2/A_{-},-1]\times F/(-1-2/A_{-},p)\sim(-1-2/A_{-},q))/G_{-}$ and push the $\mathsf{RCD}(\bar{K}_{-},N^{F}_{-}+1+N^{P}_{-})$-structure $(d_{-},\mathfrak{m}_{-})$ of $(P_{-}\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ to get an $\mathsf{RCD}(\bar{K}_{-},N^{F}_{-}+1+N^{P}_{-})$-structure on $C_{-}$, which we also denote by $(d_{-},\mathfrak{m}_{-})$ abusing notation. We also identify $(P_{+}\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ with $C_{+}:=(P_{+}\times([1,1+2/A_{+}]\times F/(1+2/A_{+},p)\sim(1+2/A_{+},q))/G_{+}$ and push the $\mathsf{RCD}(\bar{K}_{+},N^{F}_{+}+1+N^{P}_{+})$-structure $(d_{+},\mathfrak{m}_{+})$ of $(P_{+}\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ to get an $\mathsf{RCD}(\bar{K}_{+},n_{+}1+D)$-structure on $C_{+}$, which we also denote by $(d_{+},\mathfrak{m}_{+})$.
@@
[…]
We also set the measure $\mathfrak{n}_{1}:=(i_{-})_{\ast}(\mathfrak{m}_{-})+(i_{0})_{\ast}(\mathfrak{m}_{0})$, where $i_{-}\colon C_{-}\to X$ and $i_{0}\colon C_{0}\to X$ are the respective inclusions. Also observe that for $K_{1}=\min\{K_{0},\bar{K}_{+},\bar{K}_{-}\}$ and $N_{1}=\max\{N^{F}_{-}+1+N^{P}_{\pm},\dim(B)+1\}$, we have that $_{0}, CC_{-}$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces. We consider $\delta_{1}<\min\{\varepsilon,1/A_{-}\}$ as well as the open subsets $U_{-}=C_{-}\cup([-1,-1+\delta)\times B\subset X_{1}$, and $U_{0}=(-1-\delta,1]\times B$, and observe that they are an open cover of $X_{1}$ due to the fact that the metrics $d_{-}$ and $d_{0}$ are the product metrics over the preimage of $1-(2/A_{-})-\delta,-1+\delta]\subset X_{1}^{\ast}$ under the quotient map $\pi\colon X_{1}\to X_{1}^{\ast}$. Moreover, since the metric $d_{1}$ and measure $\mathfrak{m}_{1}$ are the product over $1-\delta,-1+\delta]$, we have that the spaces $(\bar{U}_{-},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{-}})$, $(\bar{U}_{0},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{0}})$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces by construction. Thus by ?THM? LABEL:T:local-to-global, we conclude that $X_{1}$ is an $\mathsf{RCD}(K_{1},N_{1})$-space.
In analogous fashion we can prove that the space
@@
We now prove ?THM? LABEL:thm.rcd_space_from_group_diagram.
Proof of ?THM? LABEL:thm.rcd_space_from_group_diagram.
We set $F_{\pm}$ to be the homogeneous Riemannian manifolds $G_{\pm}/H$ with the $G$-invariant Riemannian metrics of positive Ricci curvature when $n_{\pm}:=\mathrm{dim}(F_{\pm})>1$. Then they are $\mathsf{RCD}((n_{\pm}-1)K_{\pm},n_{\pm})$-spaces, for $n_{\pm}$ and some $K_{\pm}>0$ fixed when $n_{\pm}>1$. In the case when either $n_{+}=1$ or $n_{-}=1$, since $G_{\pm}/H$ have finite diameter $\leq D$, we can choose $K_{\pm}>0$ such that $\operatorname{diam}(F_{\pm})\leq D\leq\pi/\sqrt{K_{\pm}}$.
@@
$\Big((G\times[0,2/A_{\pm}]\times_{f_{\pm}}^{n_{\pm}}F_{\pm})/G_{\pm},(d_{G}\otimes d_{A_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes d\mathfrak{m}_{A_{\pm}})_{\ast}\Big),$
where $D=\dim G$, and $(d_{A_{\pm}},\mathfrak{m}_{A_{\pm}})$ is the distance and measure of the warped product $0,2/A_{\pm}]\times_{f_{A_{\pm}}}^{n_{\pm}}F_{\pm}$. The warping function $f_{\pm}$ is given by $f_{\pm}(t)=\min\{1,A_{\pm}t\}$. We denote by $d_{\pm}$ the metric $(d_{G}\otimes d_{A_{\pm}})^{\ast}$, and by $\mathfrak{m}_{\pm}$ the measure $(\mathfrak{m}_{G}\otimes d\mathfrak{m}_{A_{\pm}})_{\ast}$.
We showed in ?THM? LABEL:prop.near_bdry_metric_is_product_in_warped_product that over $(1/A_{\pm},2/A_{\pm}]\times F_{\pm}\subset[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}$ we have $d_{A_{\pm}}=d_{\mathbb{R}}\oplus d_{F_{\pm}}$. Since the action of $G_{\pm}$ on the $0,2/A\pm]$ factor is free, we see that the image of $G\times[1/A_{\pm},2/A_{\pm}]\times F_{\pm}$ under the quotient map $G\times([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})\to(G\times([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}))/\Delta$ is isometric to
@@
We also recall that by construction the space $B$ with the induced metric is geodesic in $(G\times[0,2/A_{\pm}]\times_{f_{\pm}}^{n_{\pm}}F_{\pm})/G_{\pm}$ with respect to the the two $\mathsf{RCD}(0,n_{\pm}+1+D)$-structures $(d_{\pm},\mathfrak{m}_{\pm})$. Moreover, over the open collars $(G((1/A_{\pm},2/A_{\pm}]\times F_{\pm}))/G_{\pm}$ the metric structures $(d_{\pm},\mathfrak{m}_{\pm})$ are the product structures $(d_{\mathbb{R}}\oplus d_{G}\oplus d_{F_{\pm}},f^{n_{\pm}}_{\pm}d\mathcal{L}\otimes d\mathfrak{m}_{G}\otimes d\mathfrak{m}_{F_{\pm}})$. From this we observe that induced metric structures on $B$ are of the form $((d_{G}\otimes d_{F_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes\mathfrak{m}_{F_{\pm}})_{\ast})$ and are given by a Riemannian metric $g_{\pm}$ on $B$: since $G\times F_{\pm}$ with $g_{G}\oplus g_{F_{\pm}}$ is a Riemannian manifold inducing $(d_{G}\oplus d_{F_{\pm}},\mathfrak{m}_{G}\oplus\mathfrak{m}_{F_{\pm}})$, and the action of $G_{\pm}$ over $G\times F_{\pm}$ is free, we obtain quotient Riemannian metrics $g_{\pm}$ on $B$ inducing $((d_{G}\oplus d_{F_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes\mathfrak{m}_{F})_{\ast})$.
We consider $1,1]\times B$, and consider $g(t)$ a path in the set of Riemannian metrics on $B$ with $g(t)=g_{-}$ over $1,-1+\varepsilon]$ and $g(t)=g_{+}$ over $1-\varepsilon,1]$. We set $C_{0}:=[-1,1]\times B$ and equip it with the metric $d_{0}$ and measure $\mathfrak{m}_{0}$ induced by the Riemannian metric $h=dt^{2}\oplus g(t)$. We observe that since $([-1,1]\times B,h)$ is a compact Riemannian manifold, then there exists $K_{0}\in\mathbb{R}$, $|K_{0}|\leq\infty$ such that $\operatorname{Ric}(h)\geq K_{0}$. This implies that $(C_{0},d_{0},\mathfrak{m}_{0})$ is an $\mathsf{RCD}(K_{0},\dim(B)+1)$-space.
We identify $0,2/A_{-}]\times_{f_{-}}^{n_{-}}F_{-}$ with $C_{-}:=(G\times([-1-2/A_{-},-1]\times F/(-1-2/A_{-},p)\sim(-1-2/A_{-},q)))/G_{-}$ and push the $\mathsf{RCD}(0,n_{-}+1+D)$-structure $(d_{-},\mathfrak{m}_{-})$ of $(G\times[0,2/A_{-}]\times_{f_{-}}^{n_{-}}F_{-})/G_{-}$ to get an $\mathsf{RCD}(0,n_{-}+1+D)$-structure on $C_{-}$, which we also denote by $(d_{-},\mathfrak{m}_{-})$ abusing notation. We also identify $G\times([0,2/A_{+}]\times_{f_{+}}^{n_{+}}F_{+})/G_{+}$ with $C_{+}:=(G\times([1,1+2/A_{+}]\times F/(1+2/A_{+},p)\sim(1+2/A_{+},q)))/G_{+}$ and push the $\mathsf{RCD}(0,n_{+}1+D)$-structure $(d_{+},\mathfrak{m}_{+})$ of $(G\times[0,2/A_{+}]\times_{f_{+}}^{n_{+}}F_{+})/G_{+}$ to get an $\mathsf{RCD}(0,n_{+}1+D)$-structure on $C_{+}$, which we also denote by $(d_{+},\mathfrak{m}_{+})$.
@@
[…]
We also set the measure $\mathfrak{n}_{1}:=(i_{-})_{\ast}(\mathfrak{m}_{-})+(i_{0})_{\ast}(\mathfrak{m}_{0})$, where $i_{-}\colon C_{-}\to X$ and $i_{0}\colon C_{0}\to X$ are the respective inclusions. Also observe that for $K_{1}=\min\{K_{0},0\}$ and $N_{1}=\max\{n_{-}+1+D,\dim(B)+1\}$, we have that $_{0}, CC_{-}$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces. We consider $\delta_{1}<\min\{\varepsilon,1/A_{-}\}$ as well as the open subsets $U_{-}=C_{-}\cup([-1,-1+\delta)\times B\subset X_{1}$, and $U_{0}=(-1-\delta,1]\times B$, and observe that they are an open cover of $X_{1}$ due to the fact that the metrics $d_{-}$ and $d_{0}$ are the product metrics over the preimage of $1-(2/A_{-})-\delta,-1+\delta]\subset X_{1}^{\ast}$ under the quotient map $\pi\colon X_{1}\to X_{1}^{\ast}$. Moreover, since the metric $d_{1}$ and measure $\mathfrak{m}_{1}$ are the product over $1-\delta,-1+\delta]$, we have that the spaces $(\bar{U}_{-},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{-}})$, $(\bar{U}_{0},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{0}})$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces by construction. Thus by ?THM? LABEL:T:local-to-global, we conclude that $X_{1}$ is an $\mathsf{RCD}(K_{1},N_{1})$-space. Observe that since the action of $G$ is transitive on $B$, we have that
$\min\{d_{-}(x,a)$ $+d_{0}((-1,a),y)\mid a\in B\}$
$=\min\{d_{-}(x,g^{-1}a)+d_{0}((-1,g^{-1}a),y)\mid g^{-1}a\in B\}$
$=\min\{d_{-}(x,\tilde{a}])+d_{0}((-1,\tilde{a}),y)\mid\tilde{w}\in B\}.$
@@
Let $x_{0}\in X$ be contained in a principal orbit, and take $k=\mathrm{dim}(G(x_{0}))$. Recall that the orbit $G(x_{0})$ is homeomorphic to the homogeneous space $G/G_{x_{0}}$. By [15, Chapter III, Section IV 54], there exists a $G$-invariant Riemannian metric $g_{0}$ on $G/G_{x_{0}}$ such that $G\subset\mathrm{Isom}(G/G_{x_{0}},g_{0})$. Hence, due to [65] and [54, Chapter II, Section 3, Theorem 3.1], we have the following dimension estimate
@@
We begin by pointing out that, if $X^{*}=\mathbf{S}^{1}$, then $X$ is homeomorphic to a $3$-manifold by item (c) of ?THM? LABEL:MC:_Homeomorphism_rigidity, and therefore the statement holds true. Thus, in the following we assume that $X^{*}=[-1,1]$.
@@
Assume first that $G=T^{3}$, and observe that all principal isotropy groups are the same for this action, since they are conjugated to each other (see [25, Theorem 4.7 and proof of Corollary 4.9]) and $G$ is abelian. Moreover, as $G$ acts by cohomogeneity one on $X$ we have a group diagram of the form:
@@
We recall from [25, Corollary 10] that $\pi^{-1}(-1,1)$ is homeomorphic to $(-1,1)\times G_{x}$, where $x\in\pi^{-1}(-1,1)$.
@@
Now from ?THM? LABEL:T:_Characterization_non-collapsed we have that $X$ is non-collapsed, and we consider $x_{\pm}\in X$ such that $x_{\pm}^{\ast}=\pm 1$. We study the tubular neighborhoods of the orbit $G(x_{\pm})$. We denote by $m_{\pm}=\mathrm{dim}(G(x_{\pm}))$, and observe that by our hypothesis $N=3$. Due to ?THM? LABEL:T:Slice-Theorem, we have that a tubular neighborhood of the orbit $G(x_{\pm})$ is homeomorphic to $G\times_{G_{\pm}}S_{\pm}$, where $S_{\pm}$ is the slice through $x_{\pm}$.